# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ 2.3 From traditional engineering to AI safety cases
The foundations of classical safety cases. In traditional engineering domains such as aerospace, rail, and nuclear power, safety cases serve as structured justifications that a system is acceptably safe for its intended use [kelly1996systematic, bloomfield2009safety]. The underlying logic is relatively stable: systems are designed to precise specifications, hazards can be systematically identified through techniques such as Fault Tree Analysis or Failure Mode and Effects Analysis, and risk is managed by design-time redundancy, physical containment, and procedural barriers. In these cases, the safety argument typically follows a deductive and hierarchical form: top-level claims (e.g., "The aircraft control system is acceptably safe") are decomposed into subclaims about components, processes, and operating conditions, supported by deterministic evidence such as design verification, test coverage results, and formal proofs.
These cases depend on predictability: the same inputs produce the same outputs, and component behaviour is well understood. Assurance rests on design sufficiency (the system behaves as specified) and evidence completeness (all hazards have been analysed). Once the system is validated against fixed requirements, safety justification largely stabilises until the next design change or certification cycle.
Why AI systems break this assurance model. AI systems, particularly generative and frontier AI models, disrupt nearly all these premises. They are not engineered line-by-line to a full behavioural specification but trained on massive, uncurated data using a low-level objective such as next-token prediction. Their performance and risks are emergent properties of both data and training scale. Consequently, their capabilities are discovered after training rather than defined before it [bommasani2021opportunities].
@@
  • Post-deployment capability discovery. The internal mechanisms of large neural networks are not interpretable in causal or mechanistic terms. Model behaviour can change qualitatively with scale, prompting configuration, or fine-tuning context. Safety arguments that rely on deterministic traceability or component-level understanding cannot be fully applied. Unexpected behaviours, such as prompt injection or role-play exploitation, emerge only after deployment-like testing [buhl2024safety].
@@
  • Model updates outside system owner control. AI applications often depend on proprietary large language models maintained by third parties, the model update may occur outside the direct control of the system owner, which breaks the assumption that the assured configuration can be frozen post-certification.. Because these models would change or update without notice, the nature and distribution of risks also evolve in ways users could not anticipate or verify. Unlike aircraft software frozen after certification, AI systems evolve continuously, through fine-tuning, retraining, data refresh, model combination, or tool access expansion. These changes invalidate static evidence. Assurance must therefore be dynamic, maintaining traceability across versions and regenerating claims and evidence as the system evolves [carlan2024dynamic].
@@
  • Capability emergence decoupled from specification. AI capabilities can emerge with scale, fine-tuning, or system-level scaffolding in ways that are not explicitly specified in requirements. As a result, safety-relevant behaviours may not be fully traceable from intended functionality to design artefacts, weakening the assumption that assurance can be completed through specification-driven decomposition alone.
@@
AI systems are not unique because they are uncertain, but because the locus and timing of uncertainty is shifted, and because assurance must operate under continual epistemic incompleteness, not just probabilistic risk. That framing avoids straw-manning traditional systems while sharpening the real distinction. Together, these properties mean that AI safety cases cannot rely on design-time determinism but must embrace evaluation-time discovery and run-time adaptation. Assurance becomes a process of continuous learning rather than static certification.
Emerging practices in AI safety cases. Several efforts illustrate how the safety-case paradigm is beginning to evolve for AI. Buhl et al.[buhl2024safety] proposes a structure that integrates traditional CAE patterns with AI-specific concepts such as model evaluation, interpretability analysis, and governance oversight. Their "frontier AI safety case" uses layered argumentation across technical, organisational, and societal domains. Cârlan et al. [carlan2024dynamic] extends this thinking into dynamic safety cases, linking claims to live metrics through "Checkable Safety Arguments" and "Safety Performance Indicators" (SPIs). The key insight is that evidence must be continually refreshed as model behaviour and deployment contexts evolve. Similarly, Clymer et al.[clymer2025example] demonstrates how a safety case can be constructed for misuse safeguards in large language models, combining threat models, red-teaming, and organisational response plans as evidence.
@@
Implications for reusable AI safety-case templates. The divergence between classical and AI safety cases motivates the need for reusable, AI-specific templates that encode appropriate logic, structure, and evidence expectations. Such templates should enable practitioners to:
@@ 3 Methodology
This study follows a Systematic Literature Review (SLR) approach based on established guidelines [kitchenham-guideline]. Figure 4 presents the high-level workflow of this research. We first developed an SLR protocol that defined the research scope, search strategy, study selection criteria, and data analysis methods. In the following subsections, we describe each of these steps in detail.
@@ 3.1 Research Questions
Population, Interventions, Comparison, Outcomes, and Context (PICOC) can properly direct the formation of research questions for a review study [shamsujjoha_2021_SLR_MDD, shamsujjoha_ICSA-2025]. The PICOC for this study is shown in Table 1, following established guidelines [petticrew-systematic-picoc, kitchenham-guideline]. In this research, the Population is the body of literature on AI safety cases, including CAE structures, safety case notations, and closely related assurance-case approaches. The Interventions of interest are the methods, processes, templates, tools, and taxonomies used to construct, analyse, or operationalise AI safety cases. We consider an Outcome successful if it contributes to a reusable safety-case framework, for example, a CAE taxonomy and associated pattern library. The relevant Context is limited to works that explicitly connect AI systems with safety case development; this excludes generic AI ethics or governance works that lack a clear safety-case linkage. Our objective in this study is to analyse how safety cases are currently constructed for AI systems and to synthesise reusable CAE structures that support requirement-aligned, auditable safety cases. Within this scope, we formulated the following key research questions:
@@
Exclude: General AI ethics and governance without an safety-case linkage; cybersecurity, privacy or access-control topics not framed as AI or safety assurance; generic risk-management checklists; non-AI safety cases unless patterns are clearly transferable to AI; non-English sources; opinion pieces without method or evidence.
RQ1
RQ2
Can we identify reusable safety case patterns for AI and demonstrate their use in a case study?
Our second RQ plan to derive generic CAE patterns for common AI safety concerns and test the framework by applying these patterns in an illustrative case study of an AI system.
@@ 3.2 Search strategy
We developed a strategy to search for papers that target AI safety and claim evidence and arguments to answer our RQs defined in Section 3.1. The goal is to find as many primary study papers as possible. Our strategy consisted of four parts: search string identification, automatic search in electronic database, snowballing using seed paper and backward snowballing using google scholar. With the assistance of the PICOC approach (Table 1), our search terms were divided into three primary concepts, as shown in Table 2. These concepts helped us to set a well-formulated search string.
@@
We used alternative spellings, abbreviations, and synonyms of search terms to increase relevant number of research papers. We used truncation and wildcard operators in finding these alternative keywords. Moreover, different supplementary key terms or phrases discovered during search iterations were added to the supportive search terms list. Our supposition is that they will collect all relevant articles. When constructing the final search query, the identified keywords, their alternatives and related terms were linked with Boolean AND (&&) and OR ($\|$) operators. The OR operator was used to concatenate the synonyms; AND to concatenate the major concepts. A generic version of the search string we used are as follows:
@@
AND
(‘claim-argument-evidence’ OR ‘CAE’ OR ‘structured argument*’
OR ‘safety justification’ OR ‘evidence-based safety’ OR ‘safety argument*’
OR ‘goal structuring notation’ OR ‘GSN’)
AND
(‘artificial intelligence’ OR AI OR ‘machine learning’
OR ‘autonomous system*’ OR ‘intelligent system*’ OR ‘software’)
@@ 3.3 Study collection and filtering
Our study filtration process is summarized in Figure 5.
Figure 5: Primary study selection process steps
We first ran the formatted search query on four major digital libraries, which returned 1,235 research papers. In parallel, we selected nine seed papers on AI safety and carried out snowballing, which returned 807 papers. The key rationale for selecting these papers as seed papers is discussed in Appendix A. We then removed 264 papers (205 from the database search results and 59 from the nine seed papers snowballing results) because they were duplicate articles, editorials or keynotes, or papers with no listed authors. After reading the title, abstract, conclusion and, when necessary, skimming the introduction, methodology and results sections, we applied Exclusion Criteria defined in Table 3. This step removed 1,588 papers in total. In the third step of filtration, we applied the Inclusion Criteria (IC) shown in Table 3 and removed a further 113 papers that did not meet our ICs.
@@
EC1 Papers about AI but not use safety or claim argument evidence as key point of concern. IC1 Studies that propose, specify, evaluate, or apply AI safety cases, claim argument evidence.
EC2 Papers discussing on AI or machine learning but not regarding safety case or CAE. IC2 Content includes sufficient methodological or technical detail to extract and explicit linkage between AI risks or regulatory requirements and the safety case structure.
EC3 Safety without relevant AI or machine learning, beyond the scope.
EC4 Papers with inadequate information to extract (Irrelevant Papers). IC3 Published or publicly available up to the review’s search date (August 2025).
EC5 Gray literature, workshop articles, posters, books, work in-progress proposals, key notes, editorial, secondary or review studies, vision papers with no concrete implementation. IC4 Full text of Conference papers, journal articles, regulatory documents, and industry credible that comply with three concepts defined in Table 2.
EC6 Discussions papers and opinion papers, as well as surveys that do not include any solution. IC5 Entire papers are written in English and use references.
EC7 Short papers less than three pages, irrelevant and low quality studies that do not contain considerable amount of information to extract. IC6 Papers available in an electronic format i.e., doc, docx, pdf, HTML, ps.
EC8 Conference or workshop papers if an extended journal version of the same paper exists.
EC9 Non-primary studies (Secondary or Tertiary Studies).
We then carried out backward snowballing using Google Scholar based on the selected 41 papers from initially selected papers from database search results. This yielded a further 1,070 papers. We followed the same process for this new set of papers and selected an additional 46 papers for inclusion that satisfied all ICs and did not meet any ECs. At each step, at least two authors independently cross-checked the selection decisions and resolved any disagreements through discussion. Finally, after quality assessment (discussed in Section 3.4), cross-checking and merging the initially selected sets, we ended up with 112 primary studies for data extraction, synthesis, and analysis. These 112 original articles are summarized in Appendix B [^1] We collected the paper list in August 2025; therefore, studies published after this date are not included in this review..
@@ 3.4 Quality assessment
We used a 1-to-5 numeric score (Very Poor, Inadequate, Moderate, Good, and Excellent) for Quality Checking (QC), applied to each study using the following eleven questions (QC1 to QC11). The questions focus on the relevance of the work in the context of AI safety cases and structured arguments, such as claim-argument-evidence.
@@
Does the study present safety cases and closely related structure framework for AI systems?
@@
Is the study’s claim, presentation and finding clear and understandable?
@@
Each paper received an average score across these criteria. We set a threshold such that if a study’s average Quality Checking (QC) score was below 2.0, we excluded it from our primary set as low quality (11 studies for selected list of 123 paper after cross check). Studies with borderline scores were discussed among the team, using the detailed extracted data to decide if they should be retained. As a result, the final analysis is based on a curated collection of high-quality publications that collectively address our research questions. This strengthens the validity of the conclusions and the proposed template framework derived from the literature. QC score for all the selected study is presented in Appendix C. In the selected primary studies, QC scores ranged from 2.73 to 4.55, with a mean of 3.37 on the 1-5 scale. Of the 112 studies in the final set, 96 ($\approx 86\%$) achieved a QC score of at least 3.0, including 22 ($\approx 20\%$) that scored above 4.0, while 16 ($\approx 14\%$) fell between 2.0 and 3.0 but were retained following team discussion of their relevance.

### v3

@@ 2.3 From traditional engineering to AI safety cases
The foundations of classical safety cases. In traditional engineering domains such as aerospace, rail, and nuclear power, safety cases serve as structured justifications that a system is acceptably safe for its intended use  [kelly1999arguing, bloomfield2009safety]. The underlying logic is relatively stable: systems are designed to precise specifications, hazards can be systematically identified through techniques such as Fault Tree Analysis or Failure Mode and Effects Analysis, and risk is managed by design-time redundancy, physical containment, and procedural barriers. In these cases, the safety argument typically follows a deductive and hierarchical form: top-level claims (e.g., "The aircraft control system is acceptably safe") are decomposed into subclaims about components, processes, and operating conditions, supported by deterministic evidence such as design verification, test coverage results, and formal proofs.
These cases depend on predictability: the same inputs produce the same outputs, and the behaviour of the components is well understood. Assurance rests on design sufficiency (the system behaves as specified) and evidence completeness (all hazards have been analysed). Once the system is validated against fixed requirements, safety justification largely stabilises until the next design change or certification cycle.
Why AI systems break this assurance model. AI systems, particularly generative and frontier AI models, disrupt nearly all these premises. They are not engineered line-by-line to a full behavioural specification, but are trained on massive, uncurated data using a low-level objective, such as next-token prediction. Their performance and risks are emergent properties of both data and training scale. Consequently, their capabilities are discovered after training rather than defined before it [bommasani2021opportunities].
@@
  • Post-deployment capability discovery. The internal mechanisms of large neural networks are not interpretable in causal or mechanistic terms. Model behaviour can change qualitatively with scale, prompting configuration or fine-tuning. Safety arguments that rely on deterministic traceability or component-level understanding cannot be fully applied. Unexpected behaviours, such as prompt injection or role-play exploitation, emerge only after deployment-like testing [buhl2024safety].
@@
  • Model updates outside system owner control. AI applications often depend on proprietary large language models maintained by third parties; model updates may occur outside the system owner’s direct control, thereby undermining the assumption that the assured configuration can be frozen post-certification. Because these models change or update without notice, the nature and distribution of risks also evolve in ways users cannot anticipate or verify. Unlike aircraft software, which is frozen after certification, AI systems evolve continuously through fine-tuning, retraining, data refresh, model combination, or expanded tool access. These changes invalidate static evidence. Assurance must therefore be dynamic, maintaining traceability across versions and regenerating claims and evidence as the system evolves [carlan2024dynamic].
@@
  • Capability emergence decoupled from specification. AI capabilities can emerge with scale, fine-tuning, or system-level scaffolding in ways that are not explicitly specified in requirements. As a result, safety-relevant behaviours may not be fully traceable from the intended functionality to the design artefacts, thereby weakening the assumption that assurance can be achieved through specification-driven decomposition alone.
@@
AI systems are not unique because they are uncertain, but because the locus and timing of uncertainty is shifted, and because assurance must operate under continual epistemic incompleteness, not just probabilistic risk. That framing avoids straw-manning traditional systems while sharpening the real distinction. Together, these properties mean that AI safety cases cannot rely on design-time determinism; instead, they must embrace evaluation-time discovery and run-time adaptation. Assurance becomes a process of continuous learning rather than static certification.
Emerging practices in AI safety cases. Several efforts illustrate how the safety-case paradigm is beginning to evolve for AI. Buhl et al. [buhl2024safety] propose a structure that integrates traditional CAE patterns with AI-specific concepts, including model evaluation, interpretability analysis, and governance oversight. Their "frontier AI safety case" uses layered argumentation across technical, organisational, and societal domains. Cârlan et al. [carlan2024dynamic] extends this thinking into dynamic safety cases, linking claims to live metrics through "Checkable Safety Arguments" and "Safety Performance Indicators" (SPIs). The key insight is that evidence must be continually refreshed as model behaviour and deployment contexts evolve. Similarly, Clymer et al. [clymer2025example] demonstrate how a safety case can be constructed for misuse safeguards in large language models, combining threat models, red-teaming, and organisational response plans as evidence.
@@
Implications for reusable AI safety-case templates. The divergence between classical and AI safety cases motivates the need for reusable, AI-specific templates that encode appropriate logic, structure, and expectations for evidence. Such templates should enable practitioners to:
@@
2.4 Positioning this study relative to existing AI safety case approaches
Prior research on AI safety cases spans multiple directions, including governance-oriented proposals, domain-specific templates, dynamic management mechanisms, and worked case sketches. However, these approaches differ in scope, level of abstraction, and reusability. To clarify the distinct contribution of this study, we compare it against representative approaches across these categories.
Table 1 positions our contribution relative to these approaches. Our contribution unifies these strands into an integrated, structured foundation for designing AI safety cases, comprising a CAE taxonomy, reusable template meta-models, and a composable pattern library that can be instantiated and systematically maintained throughout the AI lifecycle.
Table 1: Positioning of this study relative to representative AI safety-case approaches (illustrative, not exhaustive).
Approach Primary output Scope and reusability What this study adds
Safety cases for frontier AI [buhl2024safety, hilton2025scalable] Motivates safety cases for frontier AI governance; outlines practicalities of producing a frontier AI safety case Primarily high-level guidance for producing and using safety cases; does not provide an explicit and practical-level solutions Provides an integrated CAE taxonomy, reusable template, and pattern library that operationalises how to construct CAE chains across recurring AI assurance problems
Cyber inability safety-case template [goemans2024template, goemans2024template] A concrete CAE template arguing a model lacks unacceptable offensive cyber capability; decomposes claims into sub-claims supported by evaluation evidence Single risk-focused template (cyber inability); provides strong “worked CAE template” but narrow coverage across assurance problem types Generalises beyond a single template: comprehensive taxonomy of AI claim/argument/evidence types and a pattern library spanning multiple assurance scenarios, including cyber, control, evolution, thresholds
Dynamic safety cases for frontier AI [Erfan2020dynamic, Shreyas2022dynamic, Laure2025dynamic] Dynamic Safety Case Management System (DSCMS) concept for systematic, semi-automated safety-case revision over time; links claims to system state Lifecycle/update mechanism emphasised; demonstrated on a cyber-related template; less emphasis on a broad CAE taxonomy and reusable pattern design space Extends dynamic safety-case management by providing a structured CAE design space (taxonomy + reusable templates + patterns), clarifying which claim/argument/evidence types evolve and what evidence triggers updates
Sketch of an AI control safety case [korbak2025sketch, clymer2024justifysafety] Worked “sketch” of a control safety case for LLM agents; highlights key claims and evidence (e.g., control evaluations) Case-study style sketch (control measures); illustrates claim/evidence dependencies but is not a general template library Elevates such sketches into reusable patterns and templates grounded in a compositional CAE taxonomy, with explicit reasoning logics and evidence families
This study An integrated and structured CAE-based foundation for designing AI safety cases, comprising a taxonomy, reusable template meta-models, and a composable pattern library; guidance for maintaining auditability under AI evolution Designed for composability across AI system types and assurance contexts; supports reuse, traceability, and update triggers Unifies and systematises prior strands by providing: (i) an integrated CAE taxonomy, (ii) reusable template meta-models, (iii) composable AI-specific patterns, and (iv) lifecycle-aware update mechanisms grounded in explicit reasoning logics
@@ 3 Methodology
This study follows a Systematic Literature Review (SLR) approach based on established guidelines [kitchenham-guideline]. Figure 4 presents the high-level workflow of this research. We first developed an SLR protocol that defined the research scope, search strategy, study selection criteria, and data analysis methods. In the following subsections, we describe each step in detail.
@@ 3.1 Research Questions
Population, Interventions, Comparison, Outcomes, and Context (PICOC) can properly direct the formation of research questions for a review study [shamsujjoha_2021_SLR_MDD, shamsujjoha_ICSA-2025]. The PICOC for this study is shown in Table 2, following established guidelines [petticrew-systematic-picoc, kitchenham-guideline]. In this research, the Population is the body of literature on AI safety cases, including CAE structures, safety case notations, and closely related assurance-case approaches. The Interventions of interest are the methods, processes, templates, tools, and taxonomies used to construct, analyse, or operationalise AI safety cases. We consider an Outcome successful if it contributes to reusable safety-case solutions, such as a CAE taxonomy and an associated pattern library. The relevant Context is limited to works that explicitly connect AI systems with safety case development; this excludes generic AI ethics or governance works that lack a clear safety-case linkage. Our objective in this study is to analyse how safety cases are currently constructed for AI systems and to synthesise reusable CAE structures that support requirement-aligned, auditable safety cases. Within this scope, we formulated the following key research questions:
RQ1
RQ2
Can we identify reusable safety case templates and patterns for AI and demonstrate their use in a case study?
Our second RQ plans to derive generic CAE templates and patterns for common AI safety concerns and test the study by applying the templates and a selected pattern in an illustrative case study of an AI system.
@@
Exclude: General AI ethics and governance without safety-case linkage; security, privacy or access-control topics not framed as AI or safety assurance; generic risk-management checklists; non-AI safety cases unless patterns are clearly transferable to AI; non-English sources; opinion pieces without evidence/method.
@@ 3.2 Search strategy
We developed a strategy to search for papers that target AI safety and claim evidence and arguments to answer our RQs defined in Section 3.1. The goal is to find as many primary study papers as possible. Our strategy consisted of four parts: search string identification, automatic search in an electronic database, snowballing using a seed paper, and backward snowballing using Google Scholar. With the assistance of the PICOC approach (Table 2), our search terms were divided into three primary concepts, as shown in Table 3. These concepts helped us formulate a well-defined search string.
@@
We used alternative spellings, abbreviations, and synonyms of search terms to increase the number of relevant research papers. We used truncation and wildcard operators in finding these alternative keywords. Moreover, additional key terms or phrases identified during search iterations were added to the supporting search terms list. We assume they will collect all relevant articles. When constructing the final search query, the identified keywords, their alternatives, and related terms were linked with Boolean AND (&&) and OR ($\|$) operators. The OR operator was used to concatenate synonyms, and the AND operator was used to concatenate major concepts. A generic version of the search string we used is as follows:
@@
AND (‘claim-argument-evidence’ OR ‘CAE’ OR ‘structured argument*’ OR ‘safety
     justification’ OR ‘evidence-based safety’ OR ‘safety argument*’ OR
    ‘goal structuring notation’ OR ‘GSN’)
AND (‘artificial intelligence’ OR AI OR ‘machine learning’ OR ‘autonomous
     system*’ OR ‘intelligent system*’ OR ‘software’)
@@ 3.3 Study collection and filtering
Our study filtration process is summarized in Figure 5. We first ran the formatted search query on four major digital libraries, which returned 1,235 research papers. In parallel, we selected nine seed papers on AI safety and conducted snowballing, yielding 807 papers. The key rationale for selecting these papers as seed papers is discussed in Appendix A. We then removed 264 papers (205 from the database search results and 59 from the snowballing results of the nine seed papers) because they were duplicate articles, editorials, keynotes, or papers with no listed authors. After reading the title, abstract, conclusion, and, when necessary, skimming the introduction, methodology, and results sections, we applied Exclusion Criteria defined in Table 4. This step removed 1,588 papers in total. In the third step of filtration, we applied the Inclusion Criteria (IC) shown in Table 4 and removed a further 113 papers that did not meet ICs.
@@
EC1 Papers about AI but not use safety or claim argument evidence as key point of concern. IC1 Studies that propose, specify, evaluate, or apply safety cases / structured arguments for AI/ML systems or that provide foundational structured-argument constructs that are extracted to build the CAE taxonomy and template framework.
EC2 Papers discussing AI or machine learning but not regarding safety case or CAE.
EC3 Safety papers without relevant AI/ML and without safety-case/structured-argument constructs beyond the scope.
EC4 Papers with inadequate information to extract (irrelevant papers). IC2 Content includes sufficient methodological or technical detail to extract and explicit linkage between AI risks or regulatory requirements and the safety case structure.
EC5 Gray literature, workshop articles, posters, books, work-in-progress proposals, keynotes, editorials, secondary or review studies, vision papers with no concrete implementation.
EC6 Discussion papers and opinion papers, as well as surveys that do not include any solution. IC3 Published or publicly available up to the review’s search date (August 2025).
EC7 Short papers less than three pages, irrelevant and low-quality studies that do not contain a considerable amount of extractable information. IC4 Full-text conference papers, journal articles, regulatory documents, and credible industry publications that comply with the three concepts defined in Table 3.
EC8 Conference or workshop papers if an extended journal version of the same paper exists. IC5 Entire papers are written in English and use references.
EC9 Non-primary studies (secondary or tertiary studies). IC6 Papers available in electronic format (e.g., doc, docx, pdf, HTML, ps).
We then carried out backward snowballing in Google Scholar using the 41 papers selected from the database search results. This yielded a further 1,070 papers. We followed the same process for this new set of papers and selected an additional 46 papers for inclusion that satisfied all ICs and met none of the ECs. At each step, at least two authors independently cross-checked the selection decisions and resolved any disagreements through discussion. Finally, after quality assessment (discussed in Section 3.4), we cross-checked and merged the initially selected sets, resulting in 112 primary studies for data extraction, synthesis, and analysis. These 112 original articles are summarized in Appendix B [^1] We collected the paper list in August 2025; therefore, studies published after this date are not included in this review.. Among the 112 selected studies, several (e.g., [S27], [S43], [S58], [S70]) do not explicitly target AI systems but were retained because they articulate foundational principles of structured, goal-based safety argumentation. Although not recent, these works continue to underpin contemporary assurance practices through their treatment of argument decomposition and evidence traceability. Moreover, their methodological contributions are domain-independent and directly inform the CAE-based taxonomy and reusable templates developed in this paper.
Figure 5: Primary study selection process steps
@@ 3.4 Quality Assessment
We used a 1-to-5 numeric score (Very Poor, Inadequate, Moderate, Good, and Excellent) for Quality Checking (QC), applied to each study using the following eleven questions (QC1 to QC11). The questions focus on the relevance of the work to AI safety cases and to structured arguments, such as the claim-argument-evidence structure.
@@
Does the study present safety cases and a closely related structure for AI systems?
@@
Are the study’s claim, presentation, and findings clear and understandable?
@@
Each paper received an average score across these criteria. We set a threshold: if a study’s average Quality Checking (QC) score was below 2.0, we excluded it from our primary set as low-quality (11 studies from the selected list of 123 papers after cross-checking). Studies with borderline scores were discussed among the team, using the detailed extracted data to decide if they should be retained. As a result, the final analysis is based on a curated collection of high-quality publications that collectively address our research questions. This strengthens the validity of the conclusions and the proposed templates derived from the literature. QC score for all the selected studies is presented in Appendix C. In the selected primary studies, QC scores ranged from 2.73 to 4.55, with a mean of 3.37 on the 1-5 scale. Of the 112 studies in the final set, 96 ($\approx 86\%$) achieved a QC score of at least 3.0, including 22 ($\approx 20\%$) that scored above 4.0, while 16 ($\approx 14\%$) fell between 2.0 and 3.0 but were retained following team discussion of their relevance.